OpenStudy (anonymous):

a circle having a centre O has a triangle PQS inside it with the base QP is not on centre O. PT and QT are two tangents from an exterior point T (S, O and T are linear points - lie in one line ) it is given that angle ptq = 25 deg . find angle PSQ and OPQ

Directrix (directrix):

Is this how the diagram looks? See attachment.

OpenStudy (anonymous):

yes

Directrix (directrix):

Angle QTP is an angle formed by two tangents to a circle from an outside point. The formula for calculating its measure is shown on the attachment.

Directrix (directrix):

In the case of two tangents to a circle, the near arc and the far arc together form the entire circle.

Directrix (directrix):

The near arc is red; the far arc is blue in the diagram for the problem.

Directrix (directrix):

If the red arc has measure R, then the blue arc has measure 360 - R. 25 = 1/2[ (360 -R) - R ] Solve for R. @msingh

OpenStudy (anonymous):

50=360-2R 2R=360-50 2R=310 R=155

Directrix (directrix):

Correct. Notice that the near red arc and the angle formed by the tangents sums to 180. That will always be the case in this type scenario.

Directrix (directrix):

Look at angle S. It is an inscribed angle of a circle and has measure half of its intercepted arc. Angle S intercepts the red arc. What is the measure of angle S ? @msingh

OpenStudy (anonymous):

155

Directrix (directrix):

Try again. Hint: inscribed angle of a circle and has measure **half** of its intercepted arc

Directrix (directrix):

@msingh

OpenStudy (anonymous):

oopssrry 155/2

Directrix (directrix):

Tangents drawn to a circle from an outside point are congruent. That makes triangle PQT isosceles with segment TQ congruent to segment TP.

Directrix (directrix):

Looking at triangle PQT, recall that if two sides of a triangle are congruent, then the angles opposite those sides are congruent.

Directrix (directrix):

Angle TQP has the same measure as Angle TPQ. The sum of the interior angles of a triangle is 180. Angle T was given to be 25. So, what is the degree measure of angle TPQ? @msingh

OpenStudy (anonymous):

180-25=2x 155/2=x where x= angle TOP =angleTPQ

Directrix (directrix):

Yes.

Directrix (directrix):

Note on the attached diagram that segment OP is a radius of the circle. Also, recall that if a radius is drawn to a tangent to a circle at the point of tangency, the radius is perpendicular to the tangent.

Directrix (directrix):

Angle TPO is a right angle. Angle TPQ and Angle QPO form that right angle, making the sum of those measures 90. Angle TPQ is 77.5 or 155/2 so what is the measure of angle OPQ ? @msingh

OpenStudy (anonymous):

90-77.5

Directrix (directrix):

I think I made a typo on the angle. Regardless, you read my mind. Angle OPQ is 22.5 I think the problem is solved.

OpenStudy (anonymous):

k

Directrix (directrix):

Questions?

OpenStudy (anonymous):

till now , no questions but wanna say thank u 2 u @Directrix

Directrix (directrix):

Happy to help.